# Magnetic Materials for Quantum Magnonics

### v1

@@ III YIG films for quantum magnonics
In general, the broad garnet family Enke et al. (1978); Yoshimoto et al. (2019) offers promising diamagnetic substrates with lattice constants closely matching YIG. For example, yttrium scandium aluminum garnet (YSAG), used as a diamagnetic spacer between YIG and GGG, has been shown to reduce damping in YIG films at low temperatures Guo et al. (2022). A stronger candidate is yttrium scandium gallium garnet (YSGG), which can serve directly as a substrateGuo et al. (2023); Legrand et al. (2025). YIG/YSGG has caught up with YIG/GGG in terms of FMR linewidth at low temperaturesYoussef et al. (2025) (see Tab. 2) and even demonstrated robust and superior spin-wave propagation at 2 KAbrão et al. (2025). However, this improvement comes at the cost of lower saturation magnetization, which yields slower dipolar magnons, as the group velocity of magetostatical spin waves are pratically directly proportional to the magnetization saturation. In addition, for YIG/YSGG, only thin films of about 60 nm can be grown fully strained due to the large lattice mismatch, and these still do not reach the best room-temperature damping values achieved by YIG grown on GGG substratesSerha et al. (2025d).
@@
A breakthrough in the field came with the introduction of yttrium scandium gallium aluminum garnet (YSGAG) by C. Guguschev and C. Dubs et al. Guguschev et al. (2025). This diamagnetic garnet enables improved lattice matching up to zero mismatch with YIG. Figure 1 (h) illustrates the $\mathrm{Y}_{3}\mathrm{(Sc,Ga,Al)}_{2}\mathrm{(Ga,Al)}_{3}\mathrm{O}_{12}$ crystal structure, highlighting selected dodecahedral, octahedral, and tetrahedral oxygen coordination coordination polyhedra with the corresponding ions at their centers. By customizing the distribution ratio of the diamagnetic ions Sc, Ga, and Al on the octahedral and tetrahedral sites, the lattice parameter can be tuned to match that of epitaxial YIG films Guguschev et al. (2025). As shown in Fig. 2, taken from the studySerha et al. (2025d), YIG films grown by liquid-phase epitaxy (LPE) on YSGAG exhibit the same low damping at room temperature and at millikelvin temperatures that YIG on GGG achieves only at room temperature, thereby eliminating the drawbacks at cryogenic temperatures associated with paramagnetic substrates. Using YSGAG as the substrate, record-low FMR damping in YIG films were observed also at RT, with FMR linewidths as narrow as $0.1\,\mathrm{mT}$ at frequencies up to $4\,\mathrm{GHz}$ and a Gilbert damping parameter $\alpha=4.29\times 10^{-5}$.
@@ IV Outlook
Two converging breakthrough are leading to the perspective using ultra long living propagating magnons for future solid-state quantum technologies. First, diamagnetic YSGAG substrates with a compositionally tunable lattice can be matched to YIG for nanometer-thin as well as micrometer-thick films. This approach removes substrate magnetization and stray-field gradients at cryogenic temperatures while preserving the room-temperature damping performance that has made YIG/GGG films the benchmark. Second, in ultra pure YIG up to 18 µs microsecond lifetimes for short wavelength dipolar–exchange magnons were demonstrated once the thermal magnon bath is depleted, showing that long coherence is available when impurities and extrinsic scattering are suppressed. Together these results point to YIG films that support single magnon excitation, their propagation, and manipulation on chip opening new directions for solid-state quantum magnonics.
@@
A practical route to this goal is provided by nanometer-thin YIG films grown on YSGAG, whose lattice parameters can be compositionally tuned to match YIG. For submicrometer- or micrometer-thick films, an even better lattice match between the substrate and YIG is required. Such a perfectly matched, diamagnetic system would avoid the cryogenic drawbacks of paramagnetic GGG or strained substrate/YIG combinations, allowing the film to approach the intrinsic low dissipation of bulk YIG. In this regime, magnon lifetimes are governed mainly by material imperfections, making growth and processing critical: impurity levels must approach ultra-pure single-crystal quality with strict control of rare-earth contamination, while careful surface and edge preparation suppresses two-magnon scattering. At millikelvin temperatures, depletion of the thermal magnon bath allows dipolar-exchange magnons to reach lifetimes above ten microseconds. A remaining obstacle is that DESW achieve long lifetimes by being insensitive to surfaces, but this also implies the low group velocity of dipolar waves, and their lifetime is expected to decrease with the increase in velocity. One potential solution would be to use low-frequency exchange waves with wavelengths shorter than the DESW, with reduced two-magnon scattering efficiencies and high group velocities.
@@
Taken together, these elements will provide access to long-living and long-propagating single magnons, to be used as magnonic quantum information carriers. These carriers can serve as entanglement busses or gates between superconducting qubits placed on the same YSGAG substrates. Consequently, such single magnons will enable the integration of ’classical magnonics’ data processing approaches, including RF applications, Boolean and neuromorphic computing, with magnon-based quantum computing concepts on the same nanoscale chip.

### v2

@@ III YIG films for quantum magnonics
In general, the broad garnet family Enke et al. (1978); Yoshimoto et al. (2019) offers promising diamagnetic substrates with lattice constants closely matching YIG. For example, yttrium scandium aluminum garnet (YSAG), used as a diamagnetic spacer between YIG and GGG, has been shown to reduce damping in YIG films at low temperatures Guo et al. (2022). A stronger candidate is yttrium scandium gallium garnet (YSGG), which can serve directly as a substrateGuo et al. (2023); Legrand et al. (2025). YIG/YSGG has caught up with YIG/GGG in terms of FMR linewidth at low temperaturesYoussef et al. (2025) (see Tab. 2), and even demonstrated robust and superior spin-wave propagation at 2 KAbrão et al. (2025). However, this improvement comes at the cost of lower saturation magnetization, which yields slower dipolar magnons, as the group velocity of magetostatical spin waves are pratically directly proportional to the magnetization saturation. In addition, for YIG/YSGG, only thin films of about 60 nm can be grown fully strained due to the large lattice mismatch, and these still do not reach the best room-temperature damping values achieved by YIG grown on GGG substratesSerha et al. (2025a).
@@
A breakthrough in the field came with the introduction of yttrium scandium gallium aluminum garnet (YSGAG) by C. Guguschev and C. Dubs et al. Guguschev et al. (2025). This diamagnetic garnet enables improved lattice matching up to no mismatch with YIG. Figure 1 (h) illustrates the $\mathrm{Y}_{3}\mathrm{(Sc,Ga,Al)}_{2}\mathrm{(Ga,Al)}_{3}\mathrm{O}_{12}$ crystal structure, highlighting selected dodecahedral, octahedral, and tetrahedral oxygen coordination polyhedra with the corresponding ions at their centers. By customizing the distribution ratio of the diamagnetic ions Sc, Ga, and Al on the octahedral and tetrahedral sites, the lattice parameter can be tuned to match that of epitaxial YIG films Guguschev et al. (2025). As shown in Fig. 2, taken from the studySerha et al. (2025a), YIG films grown by LPE on YSGAG exhibit simultaneously the same low damping at RT and at millikelvin temperatures that YIG on GGG achieves only at RT, thereby eliminating the drawbacks at cryogenic temperatures associated with paramagnetic substrates. Using YSGAG as the substrate, record-low FMR damping in YIG films were observed also at RT, with FMR linewidths as narrow as $0.1\,\mathrm{mT}$ at frequencies up to $4\,\mathrm{GHz}$ and a Gilbert damping parameter $\alpha=4.29\times 10^{-5}$.
@@ IV Outlook
Quantum magnonics is inherently materials-driven, with ongoing efforts focused on identifying magnetic systems whose properties best match the demands of coherence, frequency, coupling strength, and scalability for a given application. Ferromagnetic metals and Heusler compounds offer large saturation magnetization, strong exchange, and excellent compatibility with nanofabrication, enabling strong coupling to microwave photons, spin currents, and superconducting circuits, but their finite electrical conductivity introduces intrinsic dissipation that presently limits magnon lifetimes to the nanosecond regime. Organic magnets provide chemically and structurally tunable systems and have demonstrated comparatively long lifetimes for their class, reaching tens of nanoseconds, yet remain constrained by disorder, interfacial effects, and the lack of systematic investigations at millikelvin temperatures. Magnetic insulators with high magnetization saturation or anisotropy, such as hexaferrites and europium chalcogenides, open access to high-frequency magnon modes and large magnon energies, offering prospects for bias-field-free operation in waveguides and high-frequency quantum magnonics, although challenges in thin-film growth, and low-temperature damping must still be addressed. Taken together, these diverse material classes highlight that future progress in quantum magnonics will rely on targeted materials optimization across multiple platforms rather than on a single universal material system.
@@
Within this diverse materials landscape, two converging breakthrough are leading to the perspective using ultra-long-living propagating magnons for future solid-state quantum technologies. First, diamagnetic YSGAG substrates with a compositionally tunable lattice can be matched to YIG for nanometer-thin as well as micrometer-thick films. This approach removes substrate magnetization and stray-field gradients at cryogenic temperatures, while preserving the room-temperature damping performance that has made YIG/GGG films the benchmark. Second, in ultra pure YIG up to 18 µs microsecond lifetimes for short wavelength dipolar–exchange magnons were demonstrated once the thermal magnon bath is depleted, showing that long coherence is available when impurities and extrinsic scattering are suppressed. Together, these results point to YIG films that support single magnon excitation, their propagation, and manipulation on-chip opening new directions for solid-state quantum magnonics.
@@
A practical route to this goal is already provided by nanometer-thin YIG films grown on YSGAG. For submicrometer- or micrometer-thick films, an even better lattice match between the substrate and YIG is required. Such a perfectly matched, diamagnetic system would avoid the cryogenic drawbacks of paramagnetic GGG or strained substrate/YIG combinations, allowing the film to approach the intrinsic low dissipation of bulk YIG. In this regime, magnon lifetimes are governed mainly by material imperfections, making growth and processing critical: impurity levels must approach ultra-pure single-crystal quality with strict control of rare-earth contamination, while careful surface and edge preparation suppresses two-magnon scattering. At millikelvin temperatures, dipolar-exchange magnons reach lifetimes above ten microseconds. A remaining obstacle is that DESWs achieve long lifetimes by being insensitive to surfaces, but this also implies the low group velocity of dipolar waves, and their lifetime is expected to decrease with the increase in velocity. One potential solution would be to use low-frequency exchange waves with wavelengths shorter than the DESW, with reduced two-magnon scattering efficiencies and high group velocities.
@@
Taken together, these elements will provide access to long-living and long-propagating single magnons that can be used as magnonic quantum information carriers. These carriers can serve as entanglement busses or gates between superconducting qubits placed on the same YSGAG substrates. Consequently, such single magnons will enable the integration of ’classical magnonics’ data processing approaches, including RF applications, Boolean and neuromorphic computing, with magnon-based quantum computing concepts on the same nanoscale chip.